Question:

3 years ago, Blue Horizon Ltd (a company listed on Singapore Exchange operating in the logistics industry), issued 50,000 (with a face value of $1,000 each) 6% coupon (payable semi-annually) 10-year bonds at a discount of 20%. Since then, Blue Horizon Ltd has generated significant profits and its credit rating has improved from B to Baa recently. As a result, its current borrowing costs have been reduced by 50 basis points. Given the explosion of e-commerce and its recent success, the board has decided to expand its business operation into regional markets like Malaysia and Vietnam. In order to do so, they plan to issue more bonds. You have been approached by the board to advise them on how best to structure the impending bond issue such that interest cost is minimised. The earnings per share (EPS) in the most recent financial year i.e. FY 2018 was $15.50. In FY 2014, it was $12. During the most recent earnings conference call, management shared that EPS is expected to grow at 3% per year, forever. Blue Horizon Ltd adopts a 60% dividend pay-out policy and currently trades at $280.

(a) Calculate the dollar amount of coupon payable on the bond every 6 months and the yield-to-maturity of the bonds 3 years ago when it was issued by Blue Horizon Ltd. (5 marks)

(b) Calculate the yield-to-maturity and the price of each bond today. (5 marks)

(c) Recommend and justify three (3) things the board can do to reduce the coupon rate of the impending bond issuance. (6 marks)

(d) Calculate Blue Horizon Ltd’s cost of equity assuming the market applies the dividend growth model. (8 marks)

(e) Suppose the yield to maturity of the bond obtained in part (b) turned out to be higher than the cost of equity computed in part (d). Interpret and discuss this finding. (6 marks)
Answer #1
(a) $ amount of coupon payable on the bond every 6 months and the yield-to-maturity of the bonds 3 years ago when it was issued by Blue Horizon Ltd.
issued 50,000 (with a face value of $1,000 each) 6% coupon (payable semi-annually) 10-year bonds at a discount of 20%.
$ amount of coupon payable on the bond every 6 months
1000*6%/2*50000=
1500000
YTM when it was issued:
Using the Present Value of a bond formula,
PV /Price of the bond=PV of its future coupons+PV of Face Value at maturity(both discounted at the YTM)
ie. PV(of bond)=(Pmt.*(1-(1+r)^-n)/r)+(FV/(1+r)^n)
where,
PV of the bond= the discounted price of the bond,ie. 50000*1000*(1-20%)=40000000
Pmt.=semi-annual coupon amt.=50000*1000*6%/2=1500000
at r= semi-annual yield to maturity-- to be found ?
n= no.of future semi-annual coupon payments till maturity ,ie.10 yrs.*2= 20
FV=Face value to be received at maturity= 50000*$1000= 50000000
Plugging in the known values, in the above formula,
40000000=(1500000*(1-(1+r)^-20)/r)+(50000000/(1+r)^20)
Solving the above equation for r, we get the semi-annual YTM as
4.54%
Annual YTM=(1+4.54%)^2-1=
9.29%
(b)YTM and the price of each bond today
As it is given that the
current borrowing costs have been reduced by 50 basis points
100 basis points=1%.
So, 50 basis points= 1%/2=0.5%,
now the effective annual YTM will be
9.29%-0.5%=
8.79%
Now, using the formula as in a. above,
we can find the current price of the bond
at this semi-annual YTM of 8.79%/2= 4.40%
for the remaining semi-annual coupon periods of (10-3)=7 yrs.*2= 14 periods
So,
Price of the Bond =(1500000*(1-(1+0.044)^-14)/0.044)+(50000000/(1+0.044)^14)
42797304
c .3 things which the board can do to reduce the coupon rate of the impending bond issuance
1.simply fixing a lower coupon rate.
2. Increasing the bond price, so that the $ coupon amount naturally becomes lower in yield.
3. Coupon rates depend on prevailing interest rates and
4.also on the issuing company's creditworthiness-so,newer companies tend to fix lower coupon amounts, while in the infant stage--so that they can meet the periodic interest expense with certainty.
(d) Blue Horizon Ltd’s cost of equity assuming the market applies the dividend growth model.
Cost of Equity(Ke)=(D1/P0)+g
Where, D1= the next dividend
P0= Current market price
g=growth rate
ie.Ke=((15.50*(1+0.03))/280)+0.03
8.70%
e. YES.
In the above case, bond Yield 8.79% > the cost of Equity., 8.70%
may be It must also have been issued on a discount, when lesser $ proceeds from issue ,increases the yield to maturity,ie. Effective rate of interest is greater than the coupon rate.